# ABCD-LINK: Annotation Bootstrapping for Cross-Document Fine-Grained Links

### v1

@@ 6.1 Setup
For this experiment, we randomly sample 20 review-paper pairs and news article pairs. REVIEWS-HE consists of short conference papers and their corresponding peer reviews, drawn from the ARR22 subset of the NLPeer dataset. We focus on short papers to avoid long-context limitations during model inference, and select the shortest available review when multiple are present. NEWS-HE comprises article pairs from the SPICED dataset (Shushkevich et al., 2024), sampled from the Politics, Sports, and Culture categories to ensure topical diversity. In both domains, we use segment documents into sentences, followed by manual corrections to improve segmentation accuracy and remove non-content elements such as social media links or boilerplate text. See Appendix H for further details on data preparation. To support the annotation process, we used the INCEpTION platform (Klie et al., 2018), which provides a side-by-side document view and allows annotators to accept or reject pre-highlighted candidate links with access to full sentence and document context; see Appendix I for further details.
@@ 6.2 Protocol
Annotators were presented with the source and target documents side-by-side, and asked to evaluate candidate links from the source document to the target using domain-specific guidelines (Appendix L). In REVIEWS-HE, up to 8 candidate targets were shown for each source sentence: top-3 from R+LLM, top-3 from Dragon+, and 2 random distractors. In NEWS-HE, which contains shorter documents, the pool was limited to 5 candidates per source sentence: top-2 from R+LLM, top-2 from Dragon+, and one random distractor. To reduce workload, we excluded source sentences that were too short (three words or fewer), as well as explicit links (e.g. Line.X, Figure.Y) which can be resolved trivially. If a candidate target is suggested by both R+LLM and Dragon+, it was only displayed once but is counted toward both during the analysis. Annotators could accept any number of target sentences as linked to a given source, incl. no matches.
6.3 Annotation study
We recruited 15 annotators via Prolific [^5] https://www.prolific.com/ and conducted a qualification task using one "gold" document pair with 30 cross-document links labeled by the study authors for each domain (Selection criteria detailed in Appendix J). Annotators were provided with annotation guidelines for each domain. We measured agreement with the gold labels using Cohen’s $\kappa$ Cohen (1960) and selected the two annotators with the highest scores: $\kappa=0.68$ for REVIEWS-HE and $\kappa=0.72$ for NEWS-HE, both indicating substantial agreement. The selected annotators then labeled the full evaluation set in two batches of 10 document pairs per domain. After the first batch, we conducted a feedback round to clarify guideline interpretations and highlight common issues, after which the annotation resumed independently. Final inter-annotator agreement across both batches was $\kappa=0.59$ for REVIEWS-HE and $\kappa=0.60$ for NEWS-HE, reflecting substantial agreement given the subjective and open-ended nature of the task. On reviews, this exceeds the reported agreement in machine-assisted annotation by Kuznetsov et al. (2022), despite our use of crowd annotators with minimal training.
@@ 6.3 Annotation study
Table 3: Agreement rate between the two annotators on candidate links, broken down by link suggestion method. Also statistics on the number of document pairs and total labeled links for each domain.
@@ 6.5 Estimating True Recall
We note that similar to other assisted labeling settings, our setup enables large-scale comparison of linking precision, but does not provide an estimate of true recall, since only candidate links were annotated. To estimate true recall, we conducted a supplementary experiment where we exhaustively label all links on a small subset of approx. 10% of REVIEWS-HE and NEWS-HE resulting in 102 links for reviews and 80 for news. We use this subset to estimate true performance of the link suggestion approaches. As Table 4 demonstrates, across both domains, R+LLM substantially improves recall and precision over retrieval-only baselines, similar to our main results. This confirms that LLMs capture links that retrieval alone misses, and validates our framework’s effectiveness in high-recall scenarios. We provide details and additional analysis in Appendix O
@@
Table 4: True recall estimation results.
@@ 7 Conclusion
High annotation cost and the diversity of link types hinder progress in cross-document linking. We introduce a domain-agnostic framework for sentence-level linking that bootstraps annotation with no labeled data upfront. By generating synthetic linked documents, we enable automatic evaluation of linking approaches and identify the best-performing ones for assisted human annotation.
@@
We believe that our work does not carry ethical risks and contributes to better technological support for many socially relevant application areas such as academic quality control, journalism, fake news detection and propaganda analysis. The participants of the synthetic data validation study in the news domain contributed on a voluntary basis; the participants of the synthetic data validation study on peer reviews conducted it as a part of their employment at the authors’ institution. The crowdworkers on Prolific have been fairly compensated with 13 Euro per hour. The NLPeer dataset has a CC BY 4.0 license, and we release our corresponding derivative dataset under the same license. The ECB+ corpus has a CC BY 3.0 license which allows us to release the corresponding derivative dataset under CC BY 4.0 license. The SPICED dataset has a CC BY 4.0 license, but the licensing for its source material (news) is unclear. To work around this, we release a script for reconstructing the dataset along with our added data. All datasets were used according to their intended research purpose.
@@
The goal of this work was to create a domain-agnostic framework for cross-document linking, and we demonstrate its effectiveness across two distinct domains. However, this generality comes with trade-offs. While the overall framework is link-type agnostic, each step (data generation, prompting, evaluation) is instantiated with domain-specific configurations (e.g., link definitions, sampling strategies, prompt templates). We view this as a modest cost compared to the human effort required for fully manual annotation, as it balances general applicability with minimal task-specific tuning. Domain-specific characteristics (e.g., subjective critique in reviews vs. factual alignment in news) can affect how well a general approach performs out-of-the-box. Thus, more specialized models or task formulations might yield stronger results in single-domain settings. Future work could explore adaptive modules or domain-specialized tuning within our general framework.
@@
Our approach to generating synthetic data relies on prompting an LLM to produce source documents conditioned on a target document. While results show this is effective, coverage could be improved through more targeted generation strategies—such as prompting for individual sentence-level links or applying multi-pass generation.
@@ Acknowledgments
Appendix A Datasets Conversion
In NEWS-ECB, for each document pair, we label a sentence pair as linked if they contain event mentions in the same cluster. While this conversion provides a useful signal, it introduces two key limitations: (1) the resulting labels are incomplete, as only coreferent events are annotated, and (2) some event mentions occur in headlines or titles, which leads to spurious links from titles to many other sentences. In REVIEWS-F1000, the authors didn’t adjudicate the annotations, so we only select examples where both expert annotators agreed on the label. This reduces the number of links we can use, but ensures we have a comparable dataset to REVIEWS-SYNTH.
Appendix B Quantitative Criteria
@@
Appendix C Rationale for Evaluation Criteria
@@
Appendix D Retrievers’ Results
Tables 7, 8, 9 and 10 provide the full retrieval results for all models for all datasets individually, broken down by cutoff values ($k\in\{1,3,5,7,10,20\}$).
Appendix E Error analysis.
@@ Appendix D Retrievers’ Results
Appendix F LLM Prompting Techniques
We created two prompt templates corresponding to the pairwise and listwise classification setups (Figures 4 and 5). Each prompt includes a system message that frames the task: determining whether a candidate target sentence from one document is related to a source sentence from another document, using full document context. In the pairwise setup, the model classifies one sentence pair at a time; in the listwise setup, it classifies a ranked list of candidate targets in a single pass. Both formats support four prompting configurations: no guidance, link description only, in-context examples only, and both. The prompts are designed to output binary classifications and return structured JSON outputs, facilitating scalable evaluation across top-$k$ retrieved candidates. For Phi-4 and Qwen2.5, we used vLLM Kwon et al. (2023) with the structured outputs function to ensure output consistency. The experiments with the open-source models were run on A100 GPUs. For GPT-4o, we used the OpenAI API with structured outputs too. For all models, we set temperate = 0.3 and top-p = 0.9.
@@ Appendix F LLM Prompting Techniques
Appendix G LLM-only Ablation
Figure 7 shows the prompt template used in the LLM-only setup, where the model classifies all sentences in the target document without prior retrieval. The prompt includes full document context and a single source sentence, along with a link description, and in-context examples, and instructs the LLM to evaluate each target sentence for its relevance. As in the other setups, we support four prompt configurations: no guidance, with in-context examples, with a link description, and with both. We used structured outputs here too, with temperate = 0.3 and top-p = 0.9
@@ Appendix G LLM-only Ablation
Appendix H Data Preparation for Human Evaluation
For REVIEWS-HE, we randomly sampled 20 paper-review pairs from the ARR22 subset of NLPeer from the shortest 100 ones to minimize long-context issues during model inference. In cases with multiple reviews, we selected the shortest available. Peer reviews were segmented into sentences using spaCy 777model: en_core_web_md (Honnibal et al., 2020) and manually corrected for sentence boundary errors. The paper texts came pre-segmented via NLPeer. For NEWS-HE, we sampled 20 article pairs from the SPICED dataset across the Politics, Sports, and Culture categories to ensure diversity. The SPICED dataset is a multi-document summarization dataset that pairs news articles of different sources that talk about the same topic or event, so the document pairs are naturally related, and suitable for our task.
Appendix I Annotation interface
@@
Appendix J Annotator Selection Criteria
@@
Appendix K Agreement trends and Annotation Feedback.
@@
After the first batch, we reviewed disagreement patterns. One annotator tended to reject links that relied on broader document context, favoring self-contained links. The other applied stricter interpretations of the guidelines, occasionally rejecting links that only partially aligned. Based on these observations, we gave targeted feedback, clarified borderline cases, and encouraged consistent application of the criteria. This led to improved agreement in the second batch ($\kappa=0.62$ for REVIEWS-HE, $0.64$ for NEWS-HE). Howver, this could also be attributed to the annotators getting used to the task, and thus understanding the guidelines better.
Appendix L Annotation Guidelines
Table 5 and 6 contain the criteria provided to the annotators to decide whether a sentence pair should be labeled as linked or not. Because the concept of a "link" can be broad, we provided domain-specific examples to clarify what qualifies as a meaningful connection in the context of peer reviews and news.
@@
Appendix M Synthetic Data Generation Prompts
To generate synthetic peer reviews and news articles, we designed detailed prompting templates that reflect realistic domain practices while introducing controlled variation. The peer review prompt (Figure 9) instructs the model to simulate a structured review grounded partially in the source text, while maintaining a natural and critical tone consistent with academic peer reviews. It emphasizes abstraction, editorial judgment, and partial grounding to avoid simple paraphrasing. Because the news articles from the WikinewsSum dataset were scrapped directly from the webpage, they contain some artefacts that are not content of the article. Thus, we first clean that using GPT-4o-mini with the prompt template shown in Figure 8. We chose GPT-4o-mini because it reliably removed social media links and scraping artifacts (based on manual checks), while being cost-effective. Next, we segment the cleaned articles using spaCy. For the choice of model used in synthetic document generation, we had three criteria: (1) strong general performance, (2) strong instruction-following (due to the specificity of our prompts), and (3) a long-context window (to handle long docs like papers). We selected DeepSeek-R1, as reasoning models showed strong prompt adherence. Importantly, we chose DeepSeek-R1 over models like OpenAI-o1 to avoid overlap with the OpenAI models used in the downstream classification task, thereby minimizing potential model-specific biases in evaluation (Liu et al., 2024). DeepSeek-R1 also offers comparable performance to OpenAI models on standard benchmarks (DeepSeek-AI et al., 2025). The generation prompt (Figure 10) guides the model to produce synthetic news article covering the same topic, but differing in structure, tone, and emphasis to simulate diverse editorial styles. Both generation prompts include format specifications and grounding strategies. While these prompts may not be the most optimal, they produce sufficiently high-quality and controllable outputs for our specific goal of in the linking task. Table 11 provides illustrative examples of generated sentences and their corresponding linked natural sentences in both domains
Appendix N Human Evaluation Results
@@
Appendix O Manual Recall Estimation Details
To estimate recall, we sampled document pairs from REVIEWS-HE and NEWS-HE. For each domain, we selected random source sentences and exhaustively searched the target document for related sentences according to the domain-specific link definition. This was repeated until we labeled approximately 10% of the total links: 102 in reviews and 80 in news. These links were used to compute precision, recall, and F1 for the top-performing R+LLM, retriever-only (Dragon+), and a random baseline. The manual annotation was done by one of the authors of this paper. The annotation of the 102 links in reviews took approx. 11 hours. The majority of the time was spent understanding the paper and the review, and ensuring that all possible target sentences were assessed. The 80 links in news took less time at approx. 5 hours due to the shorter documents, and the majority of the time was spent carefully reading the content of each sentence to judge whether it’s related or not. This annotation further highlighted the sizable effort needed for full manual annotation of cross-document links, and the need for automatic methods to speed up and scale up the process.
Synthetic Peer Review Generation Prompt This is a scrapped article from Wikinews. Due to the scrapping, it may contain sentences that are image captions and social media links. Remove such sentences, but do not change the content of the article. Output the cleaned article in JSON format, where the key is ’cleaned_article’ and the value is the cleaned article text.
@@
Figure 8: Prompt template for corrected scrapped news articles.

### v2

@@ 6.1 Setup
We randomly sample 20 review-paper pairs and news article pairs. REVIEWS-HE consists of short conference papers and their corresponding peer reviews, drawn from the ARR22 subset of the NLPeer dataset. We focus on short papers to avoid long-context limitations during model inference, and select the shortest available review when multiple are present. NEWS-HE comprises article pairs from the SPICED dataset (Shushkevich et al., 2024), sampled from the Politics, Sports, and Culture categories to ensure topical diversity. In both domains, we segment documents into sentences, followed by manual corrections to improve segmentation accuracy and remove non-content elements such as social media links or boilerplate text. See Appendix K for further details on data preparation. To support the annotation process, we used the INCEpTION platform (Klie et al., 2018), which provides a side-by-side document view and allows annotators to accept or reject pre-highlighted candidate links with access to full sentence and document context; see Appendix L for further details.
@@ 6.2 Protocol
Annotators were presented with the source and target documents side-by-side, and asked to evaluate candidate links from the source document to the target using domain-specific guidelines (Appendix O). In REVIEWS-HE, up to 8 candidate targets were shown for each source sentence: top-3 from R+LLM, top-3 from Dragon+, and 2 random distractors. In NEWS-HE, which contains shorter documents, the pool was limited to 5 candidates per source sentence: top-2 from R+LLM, top-2 from Dragon+, and one random distractor. To reduce workload, we excluded source sentences that were too short (three words or fewer), as well as explicit links (e.g. Line.X, Figure.Y) which can be resolved trivially. If a candidate target is suggested by both R+LLM and Dragon+, it was only displayed once but is counted toward both during the analysis. Annotators could accept any number of target sentences as linked to a given source, incl. no matches.
@@
Table 3: Acceptance rate between the two annotators on candidate links, broken down by link suggestion method. Also statistics on the number of document pairs and total labeled links for each domain.
6.3 Annotation study
We recruited 15 annotators via Prolific [^6] https://www.prolific.com/ and conducted a qualification task using one "gold" document pair with 30 cross-document links labeled by the study authors for each domain (Selection criteria detailed in Appendix M). Annotators were provided with annotation guidelines for each domain. We measured agreement with the gold labels using Cohen’s $\kappa$ Cohen (1960) and selected the two annotators with the highest scores: $\kappa=0.68$ for REVIEWS-HE and $\kappa=0.72$ for NEWS-HE, both indicating substantial agreement. The selected annotators then labeled the full evaluation set in two batches of 10 document pairs per domain. After the first batch, we conducted a feedback round to clarify guideline interpretations and highlight common issues, after which the annotation resumed independently. Final inter-annotator agreement across both batches was $\kappa=0.59$ for REVIEWS-HE and $\kappa=0.60$ for NEWS-HE, reflecting substantial agreement given the subjective and open-ended nature of the task. On reviews, this exceeds the reported agreement in machine-assisted annotation by Kuznetsov et al. (2022), despite our use of crowd annotators with minimal training.
@@ 6.5 Estimating True Recall
Similar to other assisted labeling settings, our setup does not provide an estimate of true recall, since only candidate links were annotated. To measure true recall, we conducted a supplementary experiment where we exhaustively labeled all links on a small subset of approx. 10% of REVIEWS-HE and NEWS-HE, resulting in 102 links for reviews and 80 for news. We use this subset to estimate the true performance of the link suggestion approaches. As Table 4 demonstrates, across both domains, R+LLM substantially improves recall and precision over retrieval-only baselines. As expected for an intersection strategy, the Both method demonstrates a drop in recall, and consequently does not outperform R+LLM on overall F1. This demonstrates that while Both can serve as a conservative high-confidence filter in annotation settings, it is not a balanced solution when considering recall. Note that in the main human study (Table 3), higher acceptance rates are due to annotating top-$k$ candidates only, whereas Table 4 measures full-corpus metrics. We provide details and additional analysis in Appendix R. In sum, LLMs capture links that retrieval alone misses, and validates our framework’s effectiveness in high-recall scenarios.
@@
Table 4: True recall estimation results under exhaustive evaluation. While “Both” achieves higher precision, its recall is substantially reduced, limiting its overall F1. Metrics are macro-averaged across source sentences.
@@ 7 Conclusion
High annotation cost and diversity of link types hinder progress in cross-document linking. We present a domain-agnostic framework for sentence-level linking that bootstraps annotation with no labeled data upfront. By generating synthetic linked documents, we enable automatic evaluation of linking approaches and identify the best-performing ones for assisted human annotation.
@@
Our work does not carry substantial additional risks and contributes to better technological support for many socially relevant application areas such as academic quality control, journalism, fake news detection and propaganda analysis. As with any AI technology, we call for additional testing and oversight if the proposed method is deployed in a sensitive domain such as law, medicine or critical infrastructure. Some general ethical risks include bias due to the use of synthetic data, lower performance in under-represented languages and domains, and dual use, for example using links between news articles and social media commentary to spot dissent. Moreover, if the generated links are used without evaluation, then incorrect links could lead to potential harms like misinformation (news domain) and wrong assumptions (peer-review domain). The participants of the synthetic data validation study in the news domain contributed on a voluntary basis; the participants of the synthetic data validation study on peer reviews conducted it as a part of their employment at the authors’ institution. The crowdworkers on Prolific have been fairly compensated with 13 Euro per hour. The NLPeer dataset has a CC BY 4.0 license, and we release our corresponding derivative dataset under the same license. The ECB+ corpus has a CC BY 3.0 license which allows us to release the corresponding derivative dataset under CC BY 4.0 license. The SPICED dataset has a CC BY 4.0 license, but the licensing for its source material (news) is unclear. To work around this, we release a script for reconstructing the dataset along with our added data. All datasets were used according to their intended research purpose.
@@
We propose a general framework for linking annotation and demonstrate its effectiveness across two distinct domains. We note that while the overall framework is agnostic to domain and link type, some steps (data generation, prompting, evaluation) need to be instantiated with domain-specific configurations (e.g., link definitions, sampling strategies, prompt templates). We view this as a modest cost compared to the human effort required for fully manual annotation, as it balances general applicability with minimal task-specific tuning. Domain-specific characteristics (e.g., subjective critique in reviews vs. factual alignment in news) can affect how well a general approach performs out-of-the-box. Thus, more specialized models or task formulations might yield stronger results in single-domain settings. Future work could explore adaptive modules or domain-specialized tuning within our general framework, as well as evaluate the applicability of our framework in further domains.
@@
Our approach to generating synthetic data relies on prompting an LLM to produce source documents conditioned on a target document. While results show this is effective, coverage could be improved through more targeted generation strategies—such as prompting for individual sentence-level links or applying multi-pass generation. That could also improve the realism of the links as discussed in Section 4.2. We also note that small human-annotated gold sets could in principle support retrieval model selection, but as shown in Section 6.5 creating them is costly and often impractical at scale. Our synthetic approach is designed for the common case where no labeled data exists upfront; when domain-specific gold data is available, it can be incorporated as a complementary signal in the evaluation stage. Finally, LLM hallucinations can lead to synthetic data that does not follow the intended goal. While we ensure data quality through human validation, targeted evaluation of LLM hallucinations in synthetic linking data generation lies beyond our scope and is left for future work.
@@ Acknowledgments
Appendix A Group-level link analysis
To assess the extent of group-level relations, we analyzed clusters of adjacent links, defined as sets of two or more links where source and/or target sentences are consecutive. Such clusters approximate larger argumentative or narrative units.
Domain Dataset #Pairs #Links %Links in Clusters Mean / Max Size
News Synthetic 346 2323 70% 3.1 / 7
Reviews Synthetic 200 2181 44% 2.4 / 7
News Human 20 296 54% 3.1 / 12
Reviews Human 20 334 21% 2.2 / 4
Table 5: Analysis of clusters of consecutive links. Group-level structures are frequent, especially in news.
The results in Table 5 confirm that group-level structures are common-particularly in news, where shorter documents promote multi-sentence alignments. Importantly, the analysis also shows that sentence-level links compose naturally into such clusters, providing atomic building blocks for broader structures. Future extensions of our framework may incorporate explicit span-level or hierarchical modeling.
Appendix B Datasets Conversion
In NEWS-ECB+, for each document pair, we label a sentence pair as linked if they contain event mentions in the same cluster. While this conversion provides a useful signal, it introduces two key limitations: (1) the resulting labels are incomplete, as only coreferent events are annotated, and (2) some event mentions occur in headlines or titles, which leads to spurious links from titles to many other sentences. In REVIEWS-F1000, the authors didn’t adjudicate the annotations, so we only select examples where both expert annotators agreed on the label. This reduces the number of links we can use, but ensures we have a comparable dataset to REVIEWS-SYNTH.
Appendix C Quantitative Criteria
@@
Appendix D Rationale for Evaluation Criteria
@@
Appendix E Retrievers’ Results
Tables 11, 12, 13 and 14 provide the full retrieval results for all models for all datasets individually, broken down by cutoff values ($k\in\{1,3,5,7,10,20\}$).
Appendix F Statistical significance testing
We conducted bootstrap resampling (10k iterations) to test whether improvements from adding LLM classification (R+LLM) over the best retriever baseline (Dragon+) are statistically significant. Results are shown in Table 6.
[…]
NEWS-SYNTH 0.74 0.84 +0.096 [0.045, 0.140] <0.001
REVIEWS-SYNTH 0.25 0.41 +0.159 [0.127, 0.188] <0.001
NEWS-ECB+ 0.55 0.57 +0.015 [-0.011, 0.042] 0.72
REVIEWS-F1000 0.59 0.59 -0.002 [-0.020, 0.032] 0.22
Table 6: Bootstrap significance test of F1 improvements from R+LLM over Dragon+.
The results confirm that improvements on the synthetic datasets are statistically significant, whereas gains on the converted datasets are not, which aligns with our earlier discussion that these datasets are less aligned with our task definition.
Appendix G Error analysis
@@
Appendix H R+LLM across different retrievers
To test whether the benefits of LLM-classification depend on the choice of base retriever, we applied the R+LLM step (Qwen2.5, listwise prompting with description + examples) to four additional retrieval models: SFR, ms_marco_MiniLM, BM25, and BGE-M3-sparse. Results are shown in Table 7.
Dataset Retriever Retriever Only R+LLM Diff
NEWS-ECB+ SFR 58.19 53.88 -4.31
ms_marco_MiniLM 54.63 54.40 -0.23
BM25 45.03 53.44 +8.41
BGE-M3-sparse 48.30 53.65 +5.35
Dragon+ 55.36 56.84 +1.48
NEWS-SYNTH SFR 70.04 78.18 +8.14
ms_marco_MiniLM 77.69 80.56 +2.87
BM25 61.78 76.63 +14.85
BGE-M3-sparse 67.56 77.61 +10.05
Dragon+ 74.38 84.02 +9.64
REVIEWS-SYNTH SFR 23.09 36.16 +13.07
ms_marco_MiniLM 27.60 37.12 +9.52
BM25 22.21 35.53 +13.32
BGE-M3-sparse 25.79 38.12 +12.33
Dragon+ 25.48 41.42 +15.94
REVIEWS-F1000 SFR 42.44 42.07 -0.37
ms_marco_MiniLM 61.03 57.38 -3.65
BM25 54.75 59.14 +4.39
BGE-M3-sparse 54.11 57.97 +3.86
Dragon+ 58.73 58.56 -0.17
Table 7: Performance of R+LLM using Qwen2.5 across different retrievers. Scores are F1 on synthetic and converted datasets.
On the synthetic datasets, the R+LLM setup yields consistently positive and similarly sized improvements on REVIEWS-SYNTH ($\approx$+9–16 F1) compared to the other synthetic dataset NEWS-SYNTH ($\approx$+8–14 F1). On converted datasets (NEWS-ECB+ and REVIEWS-F1000), improvements are marginal, which we attribute to task mismatch (Section 4). Interestingly, on REVIEWS-F1000, BM25 shows the largest positive improvement among the retrievers when combined with the LLM, also achieving the highest F1 in the R+LLM setup. This matches our explanation that BM25’s lexical matches are particularly helpful in this biomedical/technical domain due to domain-specific terminology in biology and medicine. On the other datasets BM25 starts from a lower base performance than the dense retrievers, hence the LLM has more room to correct errors, yielding larger relative gains.
Appendix I LLM Prompting Techniques
We created two prompt templates corresponding to the pairwise and listwise classification setups (Figures 5 and 6). Each prompt includes a system message that frames the task: determining whether a candidate target sentence from one document is related to a source sentence from another document, using full document context. In the pairwise setup, the model classifies one sentence pair at a time; in the listwise setup, it classifies a ranked list of candidate targets in a single pass. Both formats support four prompting configurations: no guidance, link description only, in-context examples only, and both. The prompts are designed to output binary classifications and return structured JSON outputs, facilitating scalable evaluation across top-$k$ retrieved candidates. For Phi-4 and Qwen2.5, we used vLLM Kwon et al. (2023) with the structured outputs function to ensure output consistency. The experiments with the open-source models were run on A100 GPUs. For GPT-4o, we used the OpenAI API with structured outputs too. For all models, we set temperature = 0.3 and top-p = 0.9.
@@ Appendix I LLM Prompting Techniques
Appendix J LLM-only Ablation
Figure 7 shows the prompt template used in the LLM-only setup, where the model classifies all sentences in the target document without prior retrieval. The prompt includes full document context and a single source sentence, along with a link description, and in-context examples, and instructs the LLM to evaluate each target sentence for its relevance. As in the other setups, we support four prompt configurations: no guidance, with in-context examples, with a link description, and with both. We used structured outputs here too, with temperature = 0.3 and top-p = 0.9
@@
Appendix K Data Preparation for Human Evaluation
For REVIEWS-HE, we randomly sampled 20 paper-review pairs from the ARR22 subset of NLPeer from the shortest 100 ones to minimize long-context issues during model inference. In cases with multiple reviews, we selected the shortest available. Peer reviews were segmented into sentences using spaCy 888model: en_core_web_md (Honnibal et al., 2020) and manually corrected for sentence boundary errors. The paper texts came pre-segmented via NLPeer. For NEWS-HE, we sampled 20 article pairs from the SPICED dataset across the Politics, Sports, and Culture categories to ensure diversity. The SPICED dataset is a multi-document summarization dataset that pairs news articles of different sources that talk about the same topic or event, so the document pairs are naturally related, and suitable for our task.
Appendix L Annotation interface
@@
Appendix M Annotator Selection Criteria
@@
Appendix N Agreement Trends and Annotation Feedback
@@
After the first batch, we reviewed disagreement patterns. One annotator tended to reject links that relied on broader document context, favoring self-contained links. The other applied stricter interpretations of the guidelines, occasionally rejecting links that only partially aligned. Based on these observations, we gave targeted feedback, clarified borderline cases, and encouraged consistent application of the criteria. This led to improved agreement in the second batch ($\kappa=0.62$ for REVIEWS-HE, $0.64$ for NEWS-HE). However, this could also be attributed to the annotators getting used to the task, and thus understanding the guidelines better.
Appendix O Annotation Guidelines
Table 8 and 9 contain the criteria provided to the annotators to decide whether a sentence pair should be labeled as linked or not. Because the concept of a "link" can be broad, we provided domain-specific examples to clarify what qualifies as a meaningful connection in the context of peer reviews and news.
@@
Appendix P Synthetic Data Generation Prompts
To generate synthetic peer reviews and news articles, we designed detailed prompting templates that reflect realistic domain practices while introducing controlled variation. The peer review prompt (Figure 9) instructs the model to simulate a structured review grounded partially in the source text, while maintaining a natural and critical tone consistent with academic peer reviews. It emphasizes abstraction, editorial judgment, and partial grounding to avoid simple paraphrasing. Because the news articles from the WikinewsSum dataset were scraped directly from the webpage, they contain some artefacts that are not content of the article. Thus, we first clean that using GPT-4o-mini with the prompt template shown in Figure 8. We chose GPT-4o-mini because it reliably removed social media links and scraping artifacts (based on manual checks), while being cost-effective. Next, we segment the cleaned articles using spaCy. For the choice of model used in synthetic document generation, we had three criteria: (1) strong general performance, (2) strong instruction-following (due to the specificity of our prompts), and (3) a long-context window (to handle long docs like papers). We selected DeepSeek-R1, as reasoning models showed strong prompt adherence. Importantly, we chose DeepSeek-R1 over models like OpenAI-o1 to avoid overlap with the OpenAI models used in the downstream classification task, thereby minimizing potential model-specific biases in evaluation (Liu et al., 2024). DeepSeek-R1 also offers comparable performance to OpenAI models on standard benchmarks (DeepSeek-AI et al., 2025). The generation prompt (Figure 10) guides the model to produce synthetic news article covering the same topic, but differing in structure, tone, and emphasis to simulate diverse editorial styles. Both generation prompts include format specifications and grounding strategies. While these prompts may not be the most optimal, they produce sufficiently high-quality and controllable outputs for our specific goal of in the linking task. Table 15 provides illustrative examples of generated sentences and their corresponding linked natural sentences in both domains
Appendix Q Human Evaluation Results
@@
Appendix R Manual Recall Estimation Details
To estimate recall, we sampled document pairs from REVIEWS-HE and NEWS-HE. For each domain, we selected random source sentences and exhaustively searched the target document for related sentences according to the domain-specific link definition. This was repeated until we labeled approximately 10% of the total links: 102 in reviews and 80 in news. These links were used to compute precision, recall, and F1 for the top-performing R+LLM, retriever-only (Dragon+), the “Both” intersection strategy, and a random baseline. The manual annotation was done by one of the authors of this paper.
The annotation of the 102 links in reviews took approx. 11 hours. The majority of the time was spent understanding the paper and the review, and ensuring that all possible target sentences were assessed. The 80 links in news took less time at approx. 5 hours due to the shorter documents, and the majority of the time was spent carefully reading the content of each sentence to judge whether it was related. This annotation further highlighted the sizable effort needed for full manual annotation of cross-document links, and the need for automatic methods to speed up and scale up the process.
R.1 Alternative Evaluation under Candidate-Pool Constraints
As noted in Section 6, the main human evaluation (Table 3) was conducted on top-$k$ candidate pools shown to annotators. To reconcile the difference between these acceptance rates and the exhaustive recall analysis in Table 4, we repeated the recall estimation under the same top-$k$ constraints. The results are shown in Table 10.
Recall Precision F1
NEWS
R+LLM 0.77 0.93 0.82
Retriever 0.57 0.70 0.61
Both 0.68 0.95 0.76
Random 0.02 0.05 0.03
REVIEWS
R+LLM 0.43 0.64 0.49
Retriever 0.18 0.28 0.21
Both 0.24 0.52 0.32
Random 0.00 0.00 0.00
Table 10: Recall estimation under candidate-pool constraints, mimicking the human evaluation setup. “Both” regains the high precision observed in Table 3, but at the cost of recall.
These results confirm that “Both” is best understood as a conservative high-precision filter: it produces links that are very likely to be correct, but its lower recall limits standalone utility. R+LLM remains the superior balanced approach when considering overall F1 under exhaustive conditions.
R.2 Practical Guidelines for Practitioners
Based on our experiments, our framework supports multiple operating modes depending on annotation budget and risk profile:
  • Single best method (R+LLM): Balanced precision/recall with moderate annotator load. Default choice when no strong constraints exist.
  • Intersection of top-N (“Both”): Conservative high-precision filter with reduced recall. Suitable when annotator time is very limited.
  • Union of top-N: Maximizes recall at the cost of lower precision, producing larger candidate pools. Useful when coverage is critical and annotators can handle more candidates.
These modes can be selected to match application priorities, e.g., efficiency versus coverage.
News Cleaning Prompt This is a scraped article from Wikinews. Due to the scraping, it may contain sentences that are image captions and social media links. Remove such sentences, but do not change the content of the article. Output the cleaned article in JSON format, where the key is ’cleaned_article’ and the value is the cleaned article text.
@@
Figure 8: Prompt template for corrected scraped news articles.